OpenStudy (anonymous):

What are the coordinates of Q, if the midpoint of PQ is (2, 4) and the coordinates of P are (5, 9)? A. (7, 13) B.(–3, –5) C. (–1, –1) D. (7, –5)

Nnesha (nnesha):

midpoint formula ?

Nnesha (nnesha):

\[\huge\rm (\frac{ x_1 +x_2 }{ 2 } , \frac{ y_1+y_2 }{ 2 })\] midpoint formula plug in x and y values

OpenStudy (anonymous):

x1 = 2 X2 = 5 2 + 5 = 7 7 / 2 = 3.5 __________ y1 = 4 y2 = 9 4 + 9 = 13 13 / 2 = 6.5

Nnesha (nnesha):

heheh question isn't find midpoint :P i just noticed they want us to find Q coordinate

OpenStudy (anonymous):

So what do I do ._.

OpenStudy (anonymous):

Do I need to subtract ??

Nnesha (nnesha):

 solve for x_2 we can set x values by x coordinate to find x of Q